# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v1

@@ 3 Proposed Algorithm PreDeToR
We now introduce our main algorithmic contribution, PreDeToR, which does not require access to the optimal action during training in order to improve upon the demonstrator.
3.1 Follow Demonstrator and Predict Next Reward
The key idea behind DPT is to leverage the in-context learning ability of transformers to infer the reward of each arm in a given test task. By training this in-context ability on a set of training tasks, the transformer can implicitly learn structure in the task family and exploit this structure to infer rewards without trying every single arm. Thus, in contrast to DPT and AD that output actions direction, PreDeToR outputs a scalar value reward prediction for each arm. To this effect, we append a linear layer of dimension $A$ on top of a causal GPT2 model, denoted by $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$, and use an MSE loss to train the transformer to predict the reward for each action with these outputs. Note that we use $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$ to denote a reward prediction transformer and $\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$ as the transformer that predicts a distribution over actions (as in DPT and AD ). Recall that the dataset $\mathcal{H}_{m}\sim\mathcal{D}_{\text{pre}}(\cdot|m)$ where $m\sim\mathcal{T}_{\text{pre}}$. Note that dataset consists of a history of interactions $\mathcal{H}_{m}=\{I_{m,t},r_{m,t}\}_{t=1}^{n}$. Then we pass this dataset to the transformer $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$. At every round $t$ the transformer predicts the next reward for each of the actions $a\in\mathcal{A}$ for the task $m$ based on $\mathcal{H}_{m}^{t}=\{I_{m,s},r_{m,s}\}_{s=1}^{t-1}$. This predicted reward is denoted by $\widehat{r}_{m,t+1}(a)$ for each $a\in\mathcal{A}$.
@@ 3.1 Follow Demonstrator and Predict Next Reward
Exploratory Demonstrator: Observe from the loss definition in (3) that it is calculated from the observed true reward and action from the dataset $\mathcal{H}_{m}$. In order for the transformer to learn accurate reward predictions during training, we require that the weak demonstrator is sufficiently exploratory such that it collects $\mathcal{H}_{m}$ such that $\mathcal{H}_{m}$ contains some reward $r_{m,t}$ for each action $a$. We discuss in detail the impact of the demonstrator on PreDeToR (-$\tau$) training in Section A.13.
3.2 Deploying PreDeToR for the online setting
In this section, we present the entire pseudocode of our algorithm and discuss how it is deployed for online setting. In the online setting, PreDeToR, first chooses the task $m\sim\mathcal{T}_{\text{test}}$, and the training dataset $\mathcal{H}_{m}$ is initialized empty. Then at every round $t$, PreDeToR chooses $I_{t}=\mathop{\rm arg\,max}_{a\in\mathcal{A}}\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}\left(\widehat{r}_{m,t}(a)\mid\mathcal{H}^{t}_{m}\right)$ which is the action with the highest predicted reward and $\widehat{r}_{m,t}(a)$ is the predicted reward of action $a$. Note that PreDeToR is a greedy policy and may not conduct sufficient exploration. Therefore we also introduce a soft variant, PreDeToR-$\tau$  that chooses $I_{t}\sim\mathrm{softmax}^{\tau}_{a}\left(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}\left(\mathbf{\widehat{r}}_{m,t}(a)\mid\mathcal{H}^{t}_{m}\right)\right)$. Note that PreDeToR now adds the observed reward $r_{t}(I_{t})$ at round $t$ to the dataset $\mathcal{H}_{m}$ and then uses that to act greedily for the next round $t+1$. The full pseudocode of PreDeToR in the online setting is in Algorithm 1. We discuss how PreDeToR (-$\tau$) is deployed in the offline setting in Section A.15.
@@ 3.2 Deploying PreDeToR for the online setting
16:     Add $\left\{I_{t},r_{t}\right\}$ to $\mathcal{H}_{m}^{t}$
@@
4 Empirical Study: Low Data Regime
In this section, we show that PreDeToR achieves lower regret than other in-context algorithms for the low data regime. In a low data regime, every task has a short horizon $n$ which makes this a challenging setup as there is less data available per task. Therefore any good learner must learn to leverage the structure across the tasks. This setting is quite common in recommender settings where a user interaction with the system lasts only for a few rounds (Kwon et al., 2022; Tomkins et al., 2020). We also analyze the exploration strategy of PreDeToR (-$\tau$) in this setting. We study the performance of PreDeToR in the large horizon setting in Section A.7.
@@ 4 Empirical Study: Low Data Regime
(3) DPT-greedy: This baseline is the greedy approximation of the DPT algorithm from Lee et al. (2023) which is discussed in Section 2.3. Note that we choose DPT-greedy as a representative example of similar in-context decision-making algorithms studied in Lee et al. (2023); Sinii et al. (2023); Lin et al. (2023); Ma et al. (2023); Liu et al. (2023c, a) all of which require the optimal action (or its greedy approximation). We choose DPT as a representative example of this learning procedure. The DPT-greedy uses training data $\mathcal{H}_{\text{train}}$ to estimate the optimal action for each task $m$. For each task $m$, it DPT-greedy evaluates the optimal action $\widehat{a}_{m,*}=\mathop{\rm arg\,max}_{a}\sum_{t=1}^{n}\widehat{r}_{m,t}(a)$ where, $\widehat{r}_{m,t}(a)=\sum_{t=1}^{n}\widehat{r}_{m,t}\mathbf{1}(I_{t}=a)$ is the observed cumulative reward for the action $a$ in task $m$ and $\mathbf{1}(\cdot)$ is the indicator function. Note that the optimal action $\widehat{a}_{m,*}$ is evaluated using the entire dataset $\mathcal{H}_{m}$. Finally, it implements the training procedure as discussed in Section 2.3.
@@
(5) Thomp: This baseline is the stochastic $A$-actioned bandit Thompson Sampling algorithm from Thompson (1933); Agrawal and Goyal (2012); Russo et al. (2018); Zhu and Tan (2020). We briefly describe the algorithm below: At every round $t$ and each action $a$, Thomp  samples $\gamma_{m,t}(a)\sim\mathcal{N}(\widehat{\mu}_{m,t-1}(a),\sigma^{2}/N_{m,t-1}(a))$, where $N_{m,t-1}(a)$ is the number of times the action $a$ has been selected till $t-1$, and $\widehat{\mu}_{m,t-1}(a)=\frac{\sum_{s=1}^{t-1}\widehat{r}_{m,s}\mathbf{1}(I_{s}=a)}{N_{m,t-1}(a)}$ is the empirical mean. Then the action selected at round $t$ is $I_{t}=\mathop{\rm arg\,max}_{a}\gamma_{m,t}(a)$. Observe that Thomp is not a deterministic algorithm like UCB (Auer et al., 2002). So we choose Thomp as the weak demonstrator $\pi^{w}$ because it is more exploratory than UCB and also chooses the optimal action $a_{m,*}$ sufficiently large number of times. Thomp is a weak demonstrator as it does not have access to the feature set $\mathcal{X}$ for any task $m$.
@@
(1) PreDeToR (-$\tau$) learns to exploit the latent structure of the underlying tasks even when it is trained without the optimal action $a_{m,*}$ (or its approximation) and without the action features $\mathcal{X}$.
(2) PreDeToR (-$\tau$) learns to exploit the reward correlation between tasks and is able to have an accurate reward estimate of the optimal action for each task $m$.
(3) PreDeToR (-$\tau$) outperforms its weak demonstrator even when it is trained without the optimal action $a_{m,*}$ and without the feature set $\mathcal{X}$.
@@
Experimental Result: These findings are reported in Figure 1. In Figure 1(a) we show the linear bandit setting for horizon $n=25$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=10$, and $d=2$. Further empirical setting details are stated in Section A.2. The demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. Note that for this low data regime the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. This results in higher regret. Observe that PreDeToR (-$\tau$) performs quite similarly to LinUCB and lowers regret compared to Thomp which also shows that PreDeToR is able to exploit the latent linear structure and reward correlation of the underlying tasks. AD is not able to outperform its demonstrator Thomp as it is trained to predict the next action of the demonstrator. In Figure 1(b) we show the non-linear bandit setting for horizon $n=50$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=6$, and $d=2$. The demonstrator $\pi^{w}$ is the Thomp algorithm. Again we observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy. Note that PreDeToR (-$\tau$) has lower regret than LinUCB which fails to perform well in this non-linear setting due to its algorithmic design. Finally, PreDeToR-$\tau$ performs slightly better than PreDeToR in both settings as it conducts more exploration.
@@
(b) Non-linear bandit
(c) Test action distribution
(d) Test Prediction Error
@@
Figure 1: Low-Data regime. The horizontal axis is the number of rounds. Confidence bars show one standard error.
We also show how the prediction error of the optimal action by PreDeToR is small compared to LinUCB in the linear bandit setting. In Figure 1(c) we first show how the $10$ actions are distributed in the $M_{\text{test}}=200$ test tasks. In Figure 1(c) for each bar, the frequency indicates the number of tasks where the action (shown in the x-axis) is the optimal action. Then, in Figure 1(d), we show the prediction error of PreDeToR (-$\tau$) for each task $m\in[M_{\text{test}}]$. The prediction error is calculated as $(\widehat{\mu}_{m,n,*}(a)-\mu_{m,*}(a))^{2}$ where $\widehat{\mu}_{m,n,*}(a)=\max_{a}\widehat{\bm{\theta}}_{m,n}^{\top}\mathbf{x}_{m}(a)$ is the empirical mean at the end of round $n$, and $\mu_{*,m}(a)=\max_{a}\bm{\theta}_{m,*}^{\top}\mathbf{x}_{m}(a)$ is the true mean of the optimal action in task $m$. Then we average the prediction error for the action $a\in[A]$ by the number of times the action $a$ is the optimal action in some task $m$. From the Figure 1(d), we see that for actions $\{2,3,5,6,7,10\}$, the prediction error of PreDeToR is either close or smaller than LinUCB. Note that LinUCB estimates the empirical mean directly from the test task, whereas PreDeToR has a strong prior based on the training data. So PreDeToR is able to estimate the reward of the optimal action quite well from the training dataset $\mathcal{D}_{\text{pre}}$. This shows the power of PreDeToR to go beyond the in-context decision-making setting studied in Lee et al. (2023); Lin et al. (2023); Ma et al. (2023); Sinii et al. (2023); Liu et al. (2023c) which require long horizons/trajectories and optimal action during training to learn a near-optimal policy. We discuss how exploration of PreDeToR (-$\tau$) results in low cumulative regret in Section A.11. Importantly, we show that predicting the reward has significant advantages in a low-data regime as it leads to the right notion of exploration.
5 Empirical Study: New Actions
In this section, we discuss how new actions introduced both during test time and data collection influence the performance of PreDeToR (-$\tau$) and other baselines. These experiments go beyond what is studied in Lee et al. (2023); Lin et al. (2023); Ma et al. (2023); Liu et al. (2023c) as they only study the performance in setting when $\mathcal{D}_{\text{test}}=\mathcal{D}_{\text{pre}}$. Moreover, these experiments are particularly important as they show the extent to which PreDeToR(-$\tau$) is leveraging the latent structure.

### v2

@@ 3 Proposed Algorithm PreDeToR
We now introduce our main algorithmic contribution, PreDeToR (which stands for Pre-trained Decision Transformer with Reward Estimation).
3.1 Pre-training Next Reward Prediction
The key idea behind PreDeToR is to leverage the in-context learning ability of transformers to infer the reward of each arm in a given test task. By training this in-context ability on a set of training tasks, the transformer can implicitly learn structure in the task family and exploit this structure to infer rewards without trying every single arm. Thus, in contrast to DPT and AD that output actions directly, PreDeToR outputs a scalar value reward prediction for each arm. To this effect, we append a linear layer of dimension $A$ on top of a causal GPT2 model, denoted by $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$, and use a least-squares loss to train the transformer to predict the reward for each action with these outputs. Note that we use $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$ to denote a reward prediction transformer and $\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$ as the transformer that predicts a distribution over actions (as in DPT and AD ). At every round $t$ the transformer predicts the next reward for each of the actions $a\in\mathcal{A}$ for the task $m$ based on $\mathcal{H}_{m}^{t}=\{I_{m,s},r_{m,s}\}_{s=1}^{t-1}$. This predicted reward is denoted by $\widehat{r}_{m,t+1}(a)$ for each $a\in\mathcal{A}$.
@@ 3.1 Pre-training Next Reward Prediction
Exploratory Demonstrator: Observe from the loss definition in (3) that it is calculated from the observed true reward and action from the dataset $\mathcal{H}_{m}$. In order for the transformer to learn accurate reward predictions during training, we require that the weak demonstrator is sufficiently exploratory such that it collects $\mathcal{H}_{m}$ such that $\mathcal{H}_{m}$ contains some reward $r_{m,t}$ for each action $a$. We discuss in detail the impact of the demonstrator on PreDeToR (-$\tau$) training in Section 7.
3.2 Deploying PreDeToR
At deployment time, PreDeToR learns in-context to predict the mean reward of each arm on an unseen task and acts greedily with respect to this prediction. That is, at deployment time, a new task is sampled, $m\sim\mathcal{T}_{\text{test}}$, and the dataset $\mathcal{H}_{m}^{0}$ is initialized empty. Then at every round $t$, PreDeToR chooses $I_{t}=\mathop{\rm arg\,max}_{a\in\mathcal{A}}\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}\left(\widehat{r}_{m,t}(a)\mid\mathcal{H}^{t}_{m}\right)$ which is the action with the highest predicted reward and $\widehat{r}_{m,t}(a)$ is the predicted reward of action $a$. Note that PreDeToR is a greedy policy and thus may fail to conduct sufficient exploration. To remedy this potential limitation, we also introduce a soft variant, PreDeToR-$\tau$  that chooses $I_{t}\sim\mathrm{softmax}^{\tau}_{a}\left(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}\left(\mathbf{\widehat{r}}_{m,t}(a)\mid\mathcal{H}^{t}_{m}\right)\right)$. For both PreDeToR and PreDeToR-$\tau$, the observed reward $r_{t}(I_{t})$ is added to the dataset $\mathcal{H}_{m}$ and then used to predict the reward at the next round $t+1$. The full pseudocode of using PreDeToR for online interaction is shown in Algorithm 1. In Section A.14, we discuss how PreDeToR (-$\tau$) can be deployed for offline learning.
@@ 3.2 Deploying PreDeToR
16:     Add $\left\{I_{t},r_{t}\right\}$ to $\mathcal{H}_{m}^{t}$ to form $\mathcal{H}_{m}^{t+1}$.
@@
4 Empirical Study: Non-Linear Structure
Having introduced PreDeToR, we now investigate its performance in diverse bandit settings compared to other in-context learning algorithms. In our first set of experiments, we use a bandit setting with a common non-linear structure across tasks. Ideally, a good learner would leverage the structure, however, we choose the structure such that no existing algorithms are well-suited to the non-linear structure. This setting is thus a good testbed for establishing that in-context learning can discover and exploit common structure. Moreover, each task only consists of a few rounds of interactions. This setting is quite common in recommender settings where user interaction with the system lasts only for a few rounds and has an underlying non-linear structure (Kwon et al., 2022; Tomkins et al., 2020). We show that PreDeToR achieves lower regret than other in-context algorithms for the non-linear structured bandit setting. We study the performance of PreDeToR in the large horizon setting in Section A.8.
@@ 4 Empirical Study: Non-Linear Structure
(3) DPT-greedy: This baseline is the greedy approximation of the DPT algorithm from Lee et al. (2023) which is discussed in Section 2.3. Note that we choose DPT-greedy as a representative example of similar in-context decision-making algorithms studied in Lee et al. (2023); Sinii et al. (2023); Lin et al. (2023); Ma et al. (2023); Liu et al. (2023c; a) all of which require the optimal action (or its greedy approximation). DPT-greedy estimates the optimal arm using the reward estimates for each arm during each task.
@@
(5) Thomp: This baseline is the celebrated stochastic $A$-action bandit Thompson Sampling algorithm from Thompson (1933); Agrawal & Goyal (2012); Russo et al. (2018); Zhu & Tan (2020). We choose Thomp as the weak demonstrator $\pi^{w}$ as it does not make use of arm features. Thomp is also a stochastic algorithm that induces more exploration in the demonstrations.
(6) LinUCB: (Linear Upper Confidence Bound): This baseline is the Upper Confidence Bound algorithm for the linear bandit setting that leverages the linear structure and feature of the arms to select the most promising action as well as conducting exploration. We choose LinUCB as a baseline for each test task to show the limitations of algorithms that use linear feedback structure as an underlying assumption to select actions. Note that LinUCB requires oracle access to features to select actions per task.
(7) MLinGreedy: This is the multi-task linear regression bandit algorithm proposed by Yang et al. (2021). This algorithm assumes that there is a common low-dimensional feature extractor shared between the tasks and the reward of each task linearly depends on this feature extractor. We choose MLinGreedy as a baseline to show the limitations of algorithms that use linear feedback structure across tasks as an underlying assumption to select actions. Note that MLinGreedy requires oracle access to the action features to select actions as opposed to DPT, AD, and PreDeToR.
We describe in detail the baselines Thomp, LinUCB, and MLinGreedy for interested readers in Section A.2.2.
@@
Finding 1: PreDeToR (-$\tau$) lowers regret compared to other baselines under unknown, non-linear structure. It learns to exploit the latent structure of the underlying tasks from in-context data even when it is trained without the optimal action $a_{m,*}$ (or its approximation) and without action features $\mathcal{X}$.
@@
Experimental Result: These findings are reported in Figure 1. In Figure 1(a) we show the non-linear bandit setting for horizon $n=50$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=6$, and $d=2$. The demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy. Note that for this low data regime (short horizon) the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. This results in higher regret. The PreDeToR (-$\tau$) has lower regret than LinUCB, and MLinGreedy, which fail to perform well in this non-linear setting due to their algorithmic design and linear feedback assumption. Finally, PreDeToR-$\tau$ performs slightly better than PreDeToR in both settings as it conducts more exploration.
In Figure 1(b) we show the non-linear bandit setting for horizon $n=25$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=6$, and $d=2$ where the norm of the $\bm{\theta}_{m,*}$ determines the reward of the actions which also is a non-linear function $\bm{\theta}_{m,*}$ and action features. This setting is similar to the wheel bandit setting of Riquelme et al. (2018). Again, we observe that PreDeToR has lower cumulative regret than all the other baselines.
Finally in Figure 1(c) and Figure 1(d) we show the performance of PreDeToR against other baselines in real-world datasets Movielens and Yelp. The Movielens dataset consists of more than 32 million ratings of 200,000 users and 80,000 movies (Harper & Konstan, 2015) where each entry consists of user-id, movie-id, rating, and timestamp. The Yelp dataset (Asghar, 2016) consists of ratings of 1300 business categories by 150,000 users. Each entry is summarized as user-id, business-id, rating, and timestamp. Previously structured bandit works (Deshpande & Montanari, 2012; Hong et al., 2023) directly fit a linear structure or low-rank factorization to estimate the $\bm{\theta}_{m,*}$ and simulate the ratings. However, we directly use the user-ids and movie-ids (or business-ids) to build a histogram of ratings per user and calculate the mean rating per movie (or business-id) per task. Define this as the $\{\mu_{m,a}\}_{a=1}^{A}$. This is then used to simulate the rating for $n$ horizon per movie per task where the data collection algorithm is uniform sampling. Note that this does not require estimation of user or movie features, and PreDeToR (-$\tau$) learns to exploit the latent structure of user-movie (or business) rating correlations directly from the data. From Figure 1(c) and Figure 1(d) we see that PreDeToR, and PreDeToR-$\tau$ outperform all the other baselines in these settings.
(a) Non-linear bandit
(b) Feature bandit
(c) Movielens
(d) Yelp
Figure 1: Non-linear regime. The horizontal axis is the number of rounds. Confidence bars show one standard error.
5 Empirical Study: Linear Structure and Understanding the Exploration of PreDeToR
The previous experiments were conducted in a non-linear structured setting where we are unaware of a provably near-optimal algorithm. To assess how close PreDeToR’s regret is to optimal, in this section, we consider a linear setting for which there exist well-understood algorithms (Abbasi-Yadkori et al., 2011; Lattimore & Szepesvári, 2020). Such algorithms provide a strong upper bound for PreDeToR. We summarize the key finding below:
Finding 2: PreDeToR (-$\tau$) matches the performance of the optimal algorithm LinUCB in linear bandit setting as it learns to exploit the latent structure across tasks from in-context data and without access to features.
In Figure 2 we first show the linear bandit setting for horizon $n=25$, $M_{\text{pre}}=200000$, $M_{\text{test}}=200$, $A=10$, and $d=2$. Note that the length of the context (the number of rounds) is an artifact of the transformer architecture and computational complexity. This is because the self-attention takes in as input a length-$n$ sequence of tokens of size $d$, and requires $O\left(dn^{2}\right)$ time to compute the output (Keles et al., 2023). Further empirical setting details are stated in Section A.2.
We observe from Figure 2 that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy, and AD. Note that for this low data (short horizon) regime, the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. This results in higher regret. Observe that PreDeToR (-$\tau$) performs quite similarly to LinUCB and lowers regret compared to Thomp which also shows that PreDeToR is able to exploit the latent linear structure and reward correlation of the underlying tasks. Note that LinUCB is close to the optimal algorithm for this linear bandit setting. PreDeToR outperforms AD as the main objective of AD is to match the performance of its demonstrator. In this short horizon, we see that MLinGreedy performs similarly to LinUCB.
We also show how the prediction error of the optimal action by PreDeToR is small compared to LinUCB in the linear bandit setting. In Figure 2(b) we first show how the $10$ actions are distributed in the $M_{\text{test}}=200$ test tasks. In Figure 2(b) for each bar, the frequency indicates the number of tasks where the action (shown in the x-axis) is the optimal action. Then, in Figure 2(c), we show the prediction error of PreDeToR (-$\tau$) for each task $m\in[M_{\text{test}}]$. The prediction error is calculated as $(\widehat{\mu}_{m,n,*}(a)-\mu_{m,*}(a))^{2}$ where $\widehat{\mu}_{m,n,*}(a)=\max_{a}\widehat{\bm{\theta}}_{m,n}^{\top}\mathbf{x}_{m}(a)$ is the empirical mean at the end of round $n$, and $\mu_{*,m}(a)=\max_{a}\bm{\theta}_{m,*}^{\top}\mathbf{x}_{m}(a)$ is the true mean of the optimal action in task $m$. Then we average the prediction error for the action $a\in[A]$ by the number of times the action $a$ is the optimal action in some task $m$. From the Figure 2(c), we see that for actions $\{2,3,5,6,7,10\}$, the prediction error of PreDeToR is either close or smaller than LinUCB. Note that LinUCB estimates the empirical mean directly from the test task, whereas PreDeToR has a strong prior based on the training data. So PreDeToR is able to estimate the reward of the optimal action quite well from the training dataset $\mathcal{D}_{\text{pre}}$. This shows the power of PreDeToR to go beyond the in-context decision-making setting studied in Lee et al. (2023); Lin et al. (2023); Ma et al. (2023); Sinii et al. (2023); Liu et al. (2023c) which require long horizons/trajectories and optimal action during training to learn a near-optimal policy.
@@ 5 Empirical Study: Linear Structure and Understanding the Exploration of PreDeToR
(b) Test action distribution
(c) Test Prediction Error
@@
Figure 2: Linear Expt. The horizontal axis is the number of rounds. Confidence bars show one standard error.
We now state the main finding of our analysis of exploration in the linear bandit setting:
Finding 3: The PreDeToR (-$\tau$) has an implicit two-phase exploration. In the first phase, it explores with a strong prior over the in-context training data. In the second phase, once the task data has been observed for a few rounds (in-context) it switches to task-based exploration.
We first show in Figure 3(a) the training distribution of the optimal actions. For each bar, the frequency indicates the number of tasks where the action (shown in the x-axis) is the optimal action. Then in Figure 3(b) we show how the sampling distribution of DPT-greedy, PreDeToR  and PreDeToR-$\tau$ change in the first $10$ and last $10$ rounds for all the tasks where action $5$ is optimal. To plot this graph we first sum over the individual pulls of the action taken by each algorithm over the first $10$ and last $10$ rounds. Then we average these counts over all test tasks where action $5$ is optimal. From the figure Figure 3(b) we see that PreDeToR(-$\tau$) consistently pulls the action $5$ more than DPT-greedy. It also explores other optimal actions like $\{2,3,6,7,10\}$ but discards them quickly in favor of the optimal action $5$ in these tasks. This shows that PreDeToR (-$\tau$) only considers the optimal actions seen from the training data. Once sufficient observation have been observed for the task it switches to task-based exploration and samples the optimal action more than DPT-greedy.
Finally, we plot the feasible action set considered by DPT-greedy, PreDeToR, and PreDeToR-$\tau$ in Figure 3(c). To plot this graph again we consider the test tasks where the optimal action is $5$. Then we count the number of distinct actions that are taken from round $t$ up until horizon $n$. Finally we average this over all the considered tasks where the optimal action is $5$. We call this the candidate action set considered by the algorithm. From the Figure 3(c) we see that DPT-greedy explores the least and gets stuck with few actions quickly (by round $10$). Note that the actions DPT-greedy samples are sub-optimal and so it suffers a high cumulative regret (see Figure 2). PreDeToR explore slightly more than DPT-greedy, but PreDeToR-$\tau$ explores the most.
(a) Train Optimal Action Distribution
(b) Distribution of action sampling in all test tasks where action $5$ is optimal
Figure 3: Exploration Analysis of PreDeToR(-$\tau$)
6 Empirical Study: Importance of Shared Structure and Introducing New Actions
One of our central claims is that PreDeToR (-$\tau$) internally learns and leverages the shared structure across the training and testing tasks. To validate this claim, in this section, we consider the introduction of new actions at test time that do not follow the structure of training time. These experiments are particularly important as they show the extent to which PreDeToR(-$\tau$) is leveraging the latent structure and the shared correlation between the actions and rewards.